# Random Policy Enables In-Context Reinforcement Learning within Trust Horizons

### v2

@@ 3.1 Trustworthiness of the Random Policy
Let Assumptions 1 and 2 hold. Denote by $a^{*}$ the optimal action given $s_{q}$. Consider a random policy $\pi$ as well as its Q-function $Q_{\text{MDP}}^{\pi}$. Consider the trust horizon $N>\log_{\gamma}\left(\kappa(1-\gamma)/(2B)\right)-1$. Define $\kappa=\min\limits_{s_{q}\in S}\left(Q_{\text{MDP}}^{\pi}(s_{q},a^{*})-\max\limits_{a\in A\setminus\{a^{*}\}}Q_{\text{MDP}}^{\pi}(s_{q},a)\right)$. With probability at least $1-\delta$, it holds that
$\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a^{*})\geq\max_{a\in A\setminus\{a^{*}\}}\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a),\,\forall s_{q}\in S,$ (11)
when the number of episodes $N_{\text{ep}}$ satisfies
@@
Proof. See Appendix B.3. $\square$
Theorem 2 implies that the trust horizon $N$ quantifies the trustworthiness of the decision making under the random policy $\pi$ for MDP problems. Notice that $G_{1}$ in equation 12 is monotonically decreasing with respect to the trust horizon $N$ when $N>\log_{\gamma}\left(\kappa(1-\gamma)/(2B)\right)-1$ (see Lemma 2 in Appendix B.4). Thus, with a fixed number of episodes, a larger $N$ corresponds to a higher probability (smaller $\delta$) that the average reward of the optimal action under $\pi$ exceeds that of the next-best action, indicating a more reliable decision. This aligns with the intuition that a larger $N$ corresponds to a closer approximation of the infinite-horizon MDP, where the optimal action emerges under the random policy $\pi$ (by Assumption 2).
@@
5:     Empty a step list $L_{s}$
@@
9:        Add consumed steps to $L_{s}$ (add “$N+1$” if no reward is received)
@@
In the practical implementation, we present two versions of pseudo-codes for the MDP with dense and sparse rewards. In the case of dense rewards, we randomly select a query state from the state space and execute an episode of $N$ steps. Subsequently, we choose the action that maximizes $\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a)$ across the entire action space $A$. The implementation details are summarized in Algorithm 5 (see Appendix C). Furthermore, the MDP with sparse rewards are generally more challenging to solve, as the agent does not receive feedback from the environment at each step. For any query state $s_{q}$, our goal remains to select the action that maximizes $\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a)$. However, this approach may prove ineffective if $N$ is too small, as the agent may never reach the goal, resulting in a return of 00 for all actions. To enhance the practicality in the implementation of the sparse-reward MDP, we adopt a different strategy. Specifically, for any query state $s_{q}$, we prioritize actions that can achieve the goal within $N$ steps, with the actions consuming fewer steps being preferred. If no action can accomplish the goal within $N$ steps, we sample another query state until a qualified action is identified. This action is then designated as the action label associated with the query state $s_{q}$. Details of this implementation is outlined in Algorithm 3.
Having introduced the processes for collecting the context, query state, and action label under the random policy, we can now generate the pretraining dataset by integrating the aforementioned procedures (refer to Algorithm 4). Given the pretraining dataset, the model pretraining procedure as well as the offline and online deployment for SAD are summarized in Algorithm 6 (see Appendix C).
@@
5:     Collect the query state $s_{q}$ and the action label $a_{l}$ under the environment $\tau$ and the random policy $\pi$ through Algorithm 2 for MAB (Algorithm 3 for sparse-reward MDP, Algorithm 5 for dense-reward MDP)
@@ 3.2 Performance Guarantees
In this subsection, we provide theoretical guarantees of our SAD approach. Theorems 1 and 2 imply the probability of SAD selecting the optimal action within a trust horizon. Then inspired by DPT, the trajectories generated by SAD take the same distribution as those produced by a well-specified posterior sampling (Osband et al., 2013) with a high probability. We formalize this claim in the following corollary.
@@
Proof. See Appendix B.5.
$\square$
Having established the corollary above, we next investigate the regret bound of SAD in the finite MDP setting (see details in Appendix B.6). Consider the online cumulative regret of SAD over $K$ episodes in the environment $\tau$ as $\text{Regret}_{\tau}({\mathcal{F}}_{\theta})=\sum_{k=0}^{K}V_{\tau}(\pi_{\tau}^{*})-V_{\tau}(\pi_{k})$, where $\pi_{k}(\cdot\mid s_{t})={\mathcal{F}}_{\theta}(\cdot\mid{\mathcal{C}}_{k-1},s_{t})$. Then, the regret bound of SAD is formally stated as follows.
@@
Proof. See Appendix B.7.
$\square$
@@ 4 Experiments
In this section, our empirical results on five ICRL benchmark environments (Gaussian Bandits, Bernoulli Bandits, Darkroom, Darkroom-Large, Miniworld) substantiate the efficacy of our proposed SAD method. The setup of these environments are deferred to Appendix D.
We compare our proposed SAD approach with three SOTA ICRL algorithms: AD, DPT, and DIT in the aforementioned benchmark environments (see details in Appendix D.1), and consider the DPT with optimal action labels (DPT) as the oracle upper bound of SAD. Since all these methods are FM-based, we employ the same transformer architecture (causal GPT2 model (Radford et al., 2019)) and hyperparameters (number of attention layers, number of attention heads, embedding dimensions, etc) across all experiments to ensure a fair comparison. The main hyperparameters employed in this work are summarized in Tables 1-2 (refer to Appendix D.2).
In all experiments, we employ a uniform random policy to collect context, query states, and action labels, as indicated in Algorithms 1-5. Then, we pretrain and (offline/online) deploy the FM as presented in Algorithm 6 (see Appendix C).
@@
We adhere to offline and online evaluation metrics for Bandits established in (Lee et al., 2024). In the offline evaluation, we utilize the suboptimality over different context horizon, defined by $\mu_{a^{*}}-\mu_{a}$, where $\mu_{a^{*}}$ and $\mu_{a}$ represent the mean rewards over 200 test environments of the optimal arm and the selected arm, respectively. In online evaluation, we employ cumulative regret, defined by $\sum_{t=0}^{T}(\mu_{a^{*}}-\mu_{a_{t}})$, where $a_{t}$ denotes the selected arm at time $t$. Figures 2(a) and 2(b) demonstrate that our SAD approach significantly outperforms three SOTA baselines under uniform random policy, by achieving much lower suboptimality and cumulative regret. More specifically, let us define the performance improvement of SAD over baselines in the offline evaluation by
$\frac{suboptimality_{\text{baseline}}-suboptimality_{\text{SAD}}}{suboptimality_{\text{SAD}}}.$ (15)
Likewise, the performance improvement in the online evaluation is to simply replace the suboptimality by cumulative regret. Then, SAD surpasses the best baseline, DIT, by achieving $354.0\%$ performance improvements in the offline evaluation and $273.9\%$ in the online evaluation (refer to the first row of Tables 3 and 4 in Appendix D.3). To evaluate the out-of-distribution performance of SAD compared to other baselines, we test the models pretrained by all methods on the Gaussian Bandits and assess their performance on the Bernoulli Bandits with no further fine-tuning. Figures 2(c) and 2(d) illustrate that SAD still achieves lower suboptimality and cumulative regret than all other baselines, demonstrating a more robust performance in handling out-of-distribution scenarios. More specifically, SAD surpasses the best baseline, DIT, by $289.5\%$ in the offline evaluation and $313.9\%$ in the online evaluation (refer to the second row of Tables 3 and 4 in Appendix D.3).
In the environments of Darkroom and Miniworld, we use return as the evaluation metric. Moreover, we define the performance improvement of our SAD approach over the baseline methods in both offline and online evaluations by
$\frac{Return_{\text{SAD}}-Return_{\text{baseline}}}{Return_{\text{baseline}}}.$ (16)
@@
Figure 3 demonstrates that our SAD approach significantly outperforms three SOTA baselines in the Darkroom and Darkroom-Large under uniform random policy, by achieving much higher return. In the Darkroom, SAD surpasses the best baseline, DIT, by $149.3\%$ in the offline evaluation and $41.7\%$ in the online evaluation (refer to the third row of Tables 3 and 4 in Appendix D.3). Likewise, in the Darkroom-Large, SAD outperforms the best baseline, DIT, by $266.8\%$ in the offline evaluation and $24.7\%$ in the online evaluation (refer to the fourth row of Tables 3 and 4 in Appendix D.3).
@@
Figure 4 demonstrates that our SAD approach outperforms the three SOTA baselines in the Miniworld under uniform random policy, by achieving a higher return. More specifically, SAD surpasses the best baseline, DIT, by $122.1\%$ in the offline evaluation and $21.7\%$ in the online evaluation (refer to the fifth row of Tables 3 and 4 in Appendix D.3).
Tables 3 and 4 also imply that SAD significantly outperforms all baselines on average across the five ICRL benchmark environments. In the offline evaluation, SAD exceeds the best baseline DIT by $236.3\%$ on average, the second-best DPT by $2015.9\%$, and the third-best AD by $2075.2\%$. In the online evaluation, SAD surpasses DIT by $135.2\%$, DPT by $3093.8\%$, and AD by $3208.8\%$ on average. In addition to comparing SAD with the three SOTA ICRL algorithms under the uniform random policy, we also include the empirical performance of the DPT with optimal action labels (DPT) as the oracle upper bound of SAD. We observe that SAD demonstrates performance comparable to DPT in tasks such as Gaussian Bandits, Bernoulli Bandits, DarkRoom, and DarkRoom-Large. Although Miniworld introduces challenges due to its pixel-based inputs and complex environments, SAD under the random policy still achieves approximately $50\%$ of the performance of DPT. Overall, SAD is within $18.6\%$ of the performance of DPT in the offline evaluation across the five ICRL tasks, and within $12.3\%$ in the online evaluation (see details in Tables 3 and 4; Appendix D.3).
@@ 4.1 Ablation Studies
Theorems 1 implies that the uniform random policy is probabilistically trustworthy within a horizon $N$, with monotonically increasing probability of selecting the optimal action with $N$ in the MAB problem. We substantiate this observation from the theorem by empirical evidence, as presented in Figure 5(a). Furthermore, we conduct empirical investigations into the influence of the trust horizon $N$ on the performance of the MAB problem, which considers the environments of Gaussian Bandits. As expected, a larger $N$ in the MAB problem leads to a better performance (see Figures 6(a) and 6(b)).
Figure 5: The accuracy (probability) of selecting the optimal action in the MAB and MDP problems with varying trust horizon $N$.
@@
We then shift to the MDP problem with the environment of Darkroom. Notice that in our practical algorithm for the sparse-reward MDP like Darkroom (Algorithm 3), we only utilize the state-action pairs that can reach the goal within a trust horizon. Therefore, we solely record the probability/accuracy of selecting the optimal actions on those states, as presented in Figure 5(b). It shows that the accuracy monotonically decreases with respect to the trust horizon $N$, which ideally should lead to monotonically decreasing performance as well. Nonetheless, we acknowledge that this is not the case. In particular, a large trust horizon $N$ in the MDP leads to the low accuracy of selecting the optimal action, whereas a small $N$ may induce the partially short-sighted training of FM, as Algorithm 3 solely trains on the states at most $N$ steps from the goal, instead of all states (refer to Figure 7). Our numerical results in Figures 6(c) and 6(d) validate this with the fact $N=7$ performing best, and provide the empirical evidence that either an excessively large or small trust horizon can lead to suboptimality.
Figure 7: The minimal number of steps required for a query state to reach the goal (the golden star). (a): goal in the upper-right corner. (b): goal in the middle.
@@
Figure 8: Offline and online evaluations of SAD with different transformer hyperparameters: the number of attention heads ((a) and (b)), the number of attention layers ((c) and (d)). Each hyperparameter contains four independent runs with mean and standard deviation.

### v3

@@ 4.1 Trustworthiness of the Random Policy
Let Assumptions 1 and 2 hold. Define $\kappa=\min\limits_{s_{q}\in S}\big{(}Q_{\text{MDP}}^{\pi}(s_{q},a^{*})-\max\limits_{a\in A\setminus\{a^{*}\}}Q_{\text{MDP}}^{\pi}(s_{q},a)\big{)}$. Consider the trust horizon $N>\log_{\gamma}\left(\kappa(1-\gamma)/(2B)\right)-1$. The random policy is at least $(1-\delta)$-trustworthy as in Definition 2, when the number of episodes $N_{\text{ep}}$ satisfies
@@
Proof. See Appendix A.4. $\square$
Theorem 2 implies that the trust horizon $N$ and the number of episodes $N_{\text{ep}}$ quantify the trustworthiness of the decision making under the random policy $\pi$ for MDP problems. Notice that $G_{1}$ in equation 12 is monotonically decreasing with respect to the trust horizon $N$ when $N>\log_{\gamma}\left(\kappa(1-\gamma)/(2B)\right)-1$ (see Lemma 2 in Appendix A.5). Thus, with a fixed number of episodes, a larger $N$ corresponds to a higher probability (smaller $\delta$) that the average reward of the optimal action under $\pi$ exceeds that of the next-best action, indicating a more reliable decision. This aligns with the intuition that a larger $N$ corresponds to a closer approximation of the infinite-horizon MDP, where the random policy $\pi$ selects the same optimal action as that of the optimal policy (by Assumption 2). However, it is worth noting that Theorem 2 considers the worst-case guarantee of selecting the optimal actions over all states in the state space. It is therefore could be conservative. That being said, Theorem 2 still hints to the fact that one should select the action that yields the largest average reward in practice. In the practical implementation, we randomly select a query state from the state space and execute an episode of $N$ steps for each action in the action space. When the maximal return across all actions is no less than a pre-designed return threshold $\mathfrak{R}$, we choose the action that maximizes $\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a)$ across the entire action space $A$. Otherwise, we randomly sample another query state and repeat the process of evaluating each action in a horizon $N$. The implementation details are summarized in Algorithm 3.
Algorithm 3 Collecting Query States and Action Labels under Random Policy (MDP)
1:  Require: Random policy $\pi$, state space $S$, action space $A$, environment $\tau$, trust horizon $N$, return threshold $\mathfrak{R}$
@@
5:     Empty a return list $L_{r}$
@@
8:        Run an episode of $N$ steps in $\tau$ under the random policy $\pi$
9:        Add the discounted episodic return to $L_{r}$
@@
Notably, this work focuses on the grid world navigation problems where the reward is received solely upon reaching the unique goal. In this sparse reward MDP, it is worth highlighting that the action that maximizes the discounted return is essentially the same as that reaches the goal using fewest steps. This is because that the discounted return $\sum_{t}\gamma^{t}r_{t}$ has a term $\gamma^{t}$ where $t$ is the time step. Thus, fewer steps (smaller $t$) corresponds to larger discounted return (larger $\gamma^{t}$). Therefore, for any query state $s_{q}$, we prioritize actions that can achieve the goal within $N$ steps, with the actions consuming fewer steps being preferred. If no action can accomplish the goal within $N$ steps, we sample another query state until a qualified action is identified. This action is then designated as the action label associated with the query state $s_{q}$. Details of this implementation are outlined in Algorithm 5 (see Appendix B).
Having introduced the processes for collecting the context, query state, and action label under the random policy, we can now generate the pretraining dataset by integrating the aforementioned procedures (refer to Algorithm 4). Given the pretraining dataset, the model pretraining procedure as well as the offline and online deployment for SAD are summarized in Algorithm 6 (see Appendix B).
@@
5:     Collect the query state $s_{q}$ and the action label $a_{l}$ under the environment $\tau$ and the random policy $\pi$ through Algorithm 2 for MAB (Algorithm 3 for MDP)
@@ 4.2 Performance Guarantees
This subsection establishes the performance guarantees of SAD, offering deeper insights into its efficacy.
@@
Proof. See Appendix A.6. $\square$
Notice that DPT takes the same distribution as that from a well-specified posterior sampling (Lee et al., 2024), which is widely recognized as a provably sample-efficient RL algorithm (Osband et al., 2013). However, it is crucial to highlight that our SAD approach focuses only on the pretraining dataset generation, while maintaining the same pretraining process as that of DPT. Building upon Theorems 1 and 2 implying that SAD probabilistically selects the optimal action (required by DPT), Corollary 1 substantiates that our SAD approach follows the same trajectory distribution as that of the posterior sampling with probability $(1-\delta)^{l}$.
Having established the corollary above, we next investigate the regret bound of SAD in the finite MDP setting (see details in Appendix A.7). Consider the online cumulative regret of SAD over $K$ episodes in the environment $\tau$ as $\text{Regret}_{\tau}({\mathcal{F}}_{\theta})=\sum_{k=0}^{K}V_{\tau}(\pi_{\tau}^{*})-V_{\tau}(\pi_{k})$, where $\pi_{k}(\cdot\mid s_{t})={\mathcal{F}}_{\theta}(\cdot\mid{\mathcal{C}}_{k-1},s_{t})$. Then, the regret bound of SAD is formally stated as follows.
@@
Proof. See Appendix A.8. $\square$
Analogous to Corollary 1, Corollary 2 implies that SAD achieves the same regret bound as that of DPT (equation 14) with probability $(1-\delta)^{KT}$, as SAD and DPT undergo the identical pretraining process.
@@ 5 Experiments
In this section, we substantiate the efficacy of our proposed SAD method on five ICRL benchmark problems: Gaussian Bandits, Bernoulli Bandits, Darkroom, Darkroom-Large, Miniworld, which are commonly considered in the ICRL literature (Laskin et al., 2022; Lee et al., 2024; Dong et al., 2024). All these problems are challenging to solve in-context, as the test environments differ from the pretraining environments, while the parameters of the FM remain frozen during the test.
5.1 Environmental Setup
Gaussian Bandits.
We investigate a five-armed bandit problem in which the state space $S$ consists solely of a singleton state $s_{q}$. With each arm (action) pulled, the agent receives a reward. The goal is to identify the optimal arm that can maximize the cumulative reward. We consider the reward function for each arm following a Gaussian distribution with mean $\mu_{a}$ and variance $\sigma^{2}$, i.e., $R(\cdot|s_{q},a)={\mathcal{N}}(\mu_{a},\sigma^{2})$. Each arm possesses means $\mu_{a}$ drawn from a uniform distribution $U[0,1]$ and all arms share the same variance $\sigma=0.3$. We consider the pretraining and test data to have distinct Gaussian distributions with different means.
Bernoulli Bandits.
Darkroom.
Darkroom (Laskin et al., 2022; Zintgraf et al., 2019) is a two-dimensional navigation task with discrete state and action spaces. The room consists of $7\times 7$ grids ($|S|=49$), with an unknown goal randomly placed at any of these grids. The agent can select $5$ actions: go up, go down, go left, go right, or stay. The horizon length for Darkroom is $49$, meaning the agent must reach the goal within 49 moves. The challenge of this task arises from its sparse reward structure, i.e., the agent receives a reward of 1 solely upon reaching the unique goal, and 00 otherwise. Given $7\times 7=49$ available goals, we utilize $39$ of these goals ($\sim 80\%$) for pretraining and reserve the remaining 10 ($\sim 20\%$) (unseen during pretraining) for test.
Darkroom-Large.
Miniworld.
Miniworld is a three-dimensional pixel-based navigation task. The agent is situated in a room with four differently colored boxes, one of which is the target (unknown to the agent). The agent must navigate to the target box using $25\times 25\times 3$ image observations and by selecting from 4 available actions: turn left, turn right, move forward, or stay. Similar to Darkroom, the agent receives a reward of 1 only upon approaching the unique target box, and 0 otherwise. The high-dimensional pixel inputs clearly render Miniworld a much more challenging task than Darkroom and Darkroom-Large.
5.2 Numerical Results
We consider four SOTA ICRL algorithms as our baselines in this work: AD, DIT, DPT, and DPT. Since all these methods are FM-based, we employ the same transformer architecture (causal GPT2 model (Radford et al., 2019)) and hyperparameters (number of attention layers, number of attention heads, embedding dimensions, etc) across all experiments to ensure a fair comparison. The main hyperparameters employed in this work are summarized in Tables 1-2 (refer to Appendix C.1).
In all experiments, we employ a uniform random policy to collect context, query states, and action labels, as indicated in Algorithms 1-3. Then, we pretrain the FM and deploy it with two options: online and offline. In online deployment, the pretrained model collects its own context by iteratively interacting with the test environments. In offline deployment, the pretrained model directly uses a sampled context from an offline context dataset, which is pre-collected using the random policy to interact with the test environments. The details of online/offline deployments are deferred to Algorithm 6 in Appendix B.
@@ 5.2 Numerical Results
We adhere to offline and online evaluation metrics for Bandits established in (Lee et al., 2024). In the offline evaluation, we utilize the suboptimality over different context horizon, defined by $\mu_{a^{*}}-\mu_{a}$, where $\mu_{a^{*}}$ and $\mu_{a}$ represent the mean rewards over 200 test environments of the optimal arm and the selected arm, respectively. In online evaluation, we employ cumulative regret, defined by $\sum_{t=0}^{T}(\mu_{a^{*}}-\mu_{a_{t}})$, where $a_{t}$ denotes the selected arm at time $t$. Figures 3(a) and 3(b) demonstrate that our SAD approach significantly outperforms three SOTA baselines under uniform random policy, by achieving much lower suboptimality and cumulative regret. More specifically, let us define the performance improvement of SAD over baselines in the offline evaluation by $(suboptimality_{\text{baseline}}-suboptimality_{\text{SAD}})/suboptimality_{\text{SAD}}$. Likewise, the performance improvement in the online evaluation is to simply replace the suboptimality by cumulative regret. Then, SAD surpasses the best baseline, DIT, by achieving $354.0\%$ performance improvements in the offline evaluation and $273.9\%$ in the online evaluation (refer to the first row of Tables 3 and 4 in Appendix C.2). To evaluate the out-of-distribution performance of SAD compared to other baselines, we test the models pretrained by all methods on the Gaussian Bandits and assess their performance on the Bernoulli Bandits with no further fine-tuning. Figures 3(c) and 3(d) illustrate that SAD still achieves lower suboptimality and cumulative regret than all other baselines, demonstrating a more robust performance in handling out-of-distribution scenarios. More specifically, SAD surpasses the best baseline, DIT, by $289.5\%$ in the offline evaluation and $313.9\%$ in the online evaluation (refer to the second row of Tables 3 and 4 in Appendix C.2).
In the environments of Darkroom and Miniworld, we use return as the evaluation metric. Moreover, we define the performance improvement of our SAD approach over the baseline methods in both offline and online evaluations by $(Return_{\text{SAD}}-Return_{\text{baseline}})/Return_{\text{baseline}}$.
@@
Figure 4 demonstrates that our SAD approach significantly outperforms three SOTA baselines in the Darkroom and Darkroom-Large under uniform random policy, by achieving much higher return. In the Darkroom, SAD surpasses the best baseline, DIT, by $149.3\%$ in the offline evaluation and $41.7\%$ in the online evaluation (refer to the third row of Tables 3 and 4 in Appendix C.2). Likewise, in the Darkroom-Large, SAD outperforms the best baseline, DIT, by $266.8\%$ in the offline evaluation and $24.7\%$ in the online evaluation (refer to the fourth row of Tables 3 and 4 in Appendix C.2).
@@
Figure 5 demonstrates that our SAD approach outperforms the three SOTA baselines in the Miniworld under uniform random policy, by achieving a higher return. More specifically, SAD surpasses the best baseline, DIT, by $122.1\%$ in the offline evaluation and $21.7\%$ in the online evaluation (refer to the fifth row of Tables 3 and 4 in Appendix C.2).
Tables 3 and 4 also imply that SAD significantly outperforms all baselines on average across the five ICRL benchmark environments. In the offline evaluation, SAD exceeds the best baseline DIT by $236.3\%$ on average, the second-best DPT by $2015.9\%$, and the third-best AD by $2075.2\%$. In the online evaluation, SAD surpasses DIT by $135.2\%$, DPT by $3093.8\%$, and AD by $3208.8\%$ on average. In addition to comparing SAD with the three SOTA ICRL algorithms under the uniform random policy, we also include the empirical performance of the DPT with optimal action labels (DPT) as the oracle upper bound of SAD. We observe that SAD demonstrates performance comparable to DPT in tasks such as Gaussian Bandits, Bernoulli Bandits, DarkRoom, and DarkRoom-Large. Although Miniworld introduces challenges due to its pixel-based inputs and complex environments, SAD under the random policy still achieves approximately $50\%$ of the performance of DPT. Overall, SAD is within $18.6\%$ of the performance of DPT in the offline evaluation across five ICRL tasks, and within $12.3\%$ in the online evaluation (see details in Tables 3 and 4).
@@ 5.3 Ablation Studies
Theorems 1 implies that the uniform random policy is probabilistically trustworthy within a horizon $N$, with monotonically increasing probability of selecting the optimal action with $N$ in the MAB problem. We substantiate this observation from the theorem by empirical evidence, as presented in Figure 6(a). Furthermore, we conduct empirical investigations into the influence of the trust horizon $N$ on the performance of the MAB problem, which considers the environments of Gaussian Bandits. As expected, a larger $N$ in the MAB problem leads to a better performance (see Figures 7(a) and 7(b)).
Figure 6: (a) and (b): The accuracy (probability) of selecting the optimal action in the MAB and MDP problems with varying trust horizon $N$. (c) and (d): The minimal number of steps required for a query state to reach the goal (the golden star) in the upper-right corner and in the middle.
@@
We then shift to the MDP problem with the environment of Darkroom. Notice that in our practical algorithm for the sparse-reward MDP like Darkroom (Algorithm 5), we only utilize the state-action pairs that can reach the goal within a trust horizon. Therefore, we solely record the probability/accuracy of selecting the optimal actions on those states, as presented in Figure 6(b). It shows that the accuracy monotonically decreases with respect to the trust horizon $N$, which, at first glance, may lead to monotonically decreasing performance as well. Nonetheless, we acknowledge that this is not the case. In particular, a large trust horizon $N$ in the MDP leads to the low accuracy of selecting the optimal action, whereas a small $N$ may induce the partially short-sighted training of FM, as Algorithm 5 solely trains on the states at most $N$ steps from the goal, instead of all states (refer to Figures 6(c) and 6(d)). Our numerical results in Figures 7(c) and 7(d) validate this with the fact $N=7$ performing best, and provide the empirical evidence that either an excessively large or small trust horizon can lead to suboptimality.
@@
Figure 8: Offline and online evaluations of SAD with different transformer hyperparameters: the number of attention heads ((a) and (b)); the number of attention layers ((c) and (d)). Each hyperparameter contains four independent runs with mean and standard deviation.